### v1

# Entanglement suppression and quantum scars in a three-oscillator gravitational analogue

Abstract
We study quantum correlations in a tripartite system of coupled harmonic oscillators, modeling mediator-induced interactions akin to those in quantum gravity proposals. Surprisingly, despite the system’s integrability, we identify regimes where bipartite entanglement between distant oscillators vanishes even with non-zero quantum couplings. In the Heavy Mediator Regime, entanglement is restricted to sharply localized islands in parameter space, surrounded by extended regions of suppression, analogous to quantum scars in non-integrable systems. Fidelity analysis reveal that these low-entanglement states exhibit dynamical stability and spectral signatures reminiscent of quantum scars in non-integrable systems, suggesting a continuous-variable analogue of scarring tied to hidden phase-space symmetries. The Light Mediator Regime, by contrast, displays smooth entanglement generation. These results challenge entanglement-based tests of quantum gravity: the absence of entanglement need not imply classical mediation but may instead reflect a mediator dynamically constrained to a quantum subspace—akin to gravitational memory or decoherence-free subspaces. Our findings emphasize the necessity of probing the mediator’s dynamical regime to unambiguously diagnose quantum gravity.
I Introduction
The Bose-Marletto-Vedral (BMV) proposal [1, 2] has emerged as a promising avenue to experimentally probe the quantum nature of gravity. It proposes that two mesoscopic masses, each in a spatial superposition, would become entangled through their gravitational interaction if gravity is quantum, a signature absent in a classical gravitational field [3, 4, 5]. While conceptually elegant, the BMV protocol faces formidable experimental challenges, such as maintaining coherence and detecting extremely weak gravitational entanglement [6, 7, 8, 9, 10, 11]. More fundamentally, theoretical ambiguities persist, with studies questioning whether entanglement is an unambiguous signature of quantum gravity, as classical systems can mimic statistical correlations or Bell-local correlations under certain conditions, [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. These concerns highlight the need for robust observables to discriminate between classical and quantum gravitational interactions.
In this work, we present a simplified yet insightful model of tripartite coupled harmonic oscillators (CHOs) to investigate the limitations of using entanglement as a definitive witness for quantumness of interactions, particularly in the context of BMV proposal. We demonstrate that for specific parameter regimes, this integrable system exhibits vanishing logarithmic negativity ($\mathcal{E}_{\mathcal{N}}$[24] (and quantum discord ($\mathcal{D}$[25]), despite the presence of non-zero interactions and quantum states. This behavior stems from the delicate balance of coupling strengths and oscillator frequencies, showing that vanishing entanglement or discord does not preclude quantum nature of the underlying interactions or the non-classicality of the system’s state. Our model offers a tractable framework to identify potential pitfalls in entanglement-based tests of quantum gravity.
The phenomenon of vanishing entanglement in our integrable three CHOs model bears a sriking resemblance to quantum many-body scar-states in certain non-integrable systems [26, 27, 28]. Quantum scars are specific eigenstates that defy the eigenstate thermalization hypothesis [29], exhibiting unexpectedly low entanglement and persisting quantum coherences in otherwise chaotic systems. While scars arise due to fine-tuned parameters or symmetries in non-integrable models [30], our integrable system demonstrates that such low-entanglement states can also emerge from a small set of Gaussian states in the Hilbert space. Further fidelity ($\mathcal{F}$[31] of the evolved state with the initial state remains remarkably close to unity over long times for such states  [^1] While traditionally underexplored in the literature, recent studies have shown an interest in this aspect [55].. This coexistence of dynamically localized states points towards selective resonance effects or subtle hidden symmetries. Such mechanisms could complicate entanglement-based tests of quantum gravity.
Our results have implications extending beyond the BMV proposal, underscoring the necessity of developing alternative observables and refined criteria to discriminate between classical and quantum interactions. Crucially, we demonstrate that entanglement alone cannot serve as a definitive marker of quantumness — particularly in complex systems where quantum correlations may vanish despite non-trivial dynamics. Our analysis suggests that robust tests of quantum gravity must include a broader suite of quantum signatures, including non-classical correlations beyond entanglement and interaction dynamics. This work serves as a cautionary note against overreliance on simplistic quantum witnesses, highlighting the need for rigorous theoretical frameworks to interpret experimental tests of quantum gravity.
II The model and set up
To explore these questions, we analyze a simple tripartite system of quantum CHOs, designed to mimic aspects of the BMV setup as shown in Fig. (1). We investigate the evolution of $\mathcal{E}_{\mathcal{N}}$ [24] and $\mathcal{D}$ [25], as functions of system parameters and initial conditions. The system consists of three CHOs, labeled A, B, and C. Oscillator B acts as a local mediator, connecting A and C, which do not interact directly. For simplicity, oscillators A and C are assumed identical.
Figure 1: Schematic diagram of the tripartite HO system. Oscillator B mediates interaction with A and C. Interactions include self-squeezing $(\lambda_{A},\lambda_{B})$ and inter-oscillator couplings $(g_{A},g_{B})$, like parametric squeezing and mode mixing.
The Hamiltonian incorporates three types of interactions: self-squeezing, parametric squeezing, and mode mixing. The interaction between the mediator and the other oscillators resembles those in the Caldeira-Leggett model, coupling subsystems via terms like $x_{A}x_{B}$ or $x_{B}x_{C}$ [33]. These inter-oscillator couplings decompose into two qualitatively distinct components: mode mixing, which conserves particle number (e.g., $a^{\dagger}b+ab^{\dagger}$), and parametric squeezing, which does not (e.g., $ab+a^{\dagger}b^{\dagger}$[34]. This decomposition facilitates non-trivial energy redistribution and building of correlations across subsystems. For instance, correlated two-mode processes can lead to simultaneous excitation loss from modes A and B, while mode C gains energy through indirect coupling. Although the total energy of the system is conserved, these interaction channels enable rich dynamics of information and entanglement flow among the oscillators.
We include self-squeezing terms in the Hamiltonian, similar to those observed in quantum optics for light propagating through non-linear media [35, 36]. Their inclusion is motivated by the non-linear nature of gravity, where even in the linear limit, phenomena like gravitational memory exist [37, 38, 39]. Thus, considering self-squeezing is crucial for a model aiming to shed light on gravitational interactions. This framework allows us to explore scenarios where strong correlations develop between subsystem pairs, enabling us to study the conditions under which entanglement between A and C is significant or negligible. Specifically, we’re interested in two distinct physical regimes characterized by the relative strength of the self-squeezing parameter $\lambda$ compared to the oscillator frequencies $\omega$:
  1. 1.
    The heavy mediator regime (HMR), where the self-squeezing strength of the mediator oscillator B approaches its natural frequency, i.e., $\lambda_{B}\to\omega_{B}$.
  2. 2.
    The light mediator regime (LMR), where the self-squeezing strengths of the peripheral oscillators A and C approach their natural frequencies $\lambda_{A}\to\omega_{A}$.
As we show explicitly, these two regimes give rise to qualitatively different entanglement and correlation dynamics. The total Hamiltonian (in units of $\hbar$) is
$\hat{H}=\hat{H}^{(0)}_{A}+\hat{H}^{(0)}_{B}+\hat{H}^{(0)}_{C}+\hat{H}_{AB}+\hat{H}_{BC}$ (1)
where the superscript ${(0)}$ denote the free Hamiltonians of the individual HOs along with self squeezing interaction strength ($\lambda_{A},\lambda_{B}$), while the last two terms describe the interactions strengths $(g_{A},g_{B})$:
$\hat{H}^{(0)}_{A}$ $=\omega_{A}\left(a^{{\dagger}}a+\mathbb{I}/2\right)+\lambda_{A}\left(a^{2}+a^{{\dagger}2}\right)/2$ (2a)
$\hat{H}^{(0)}_{B}$ $=\omega_{B}\left(b^{{\dagger}}b+\mathbb{I}/2\right)+\lambda_{B}\left(b^{2}+b^{{\dagger}2}\right)/2$ (2b)
$\hat{H}^{(0)}_{C}$ $=\omega_{A}\left(c^{{\dagger}}c+\mathbb{I}/2\right)+\lambda_{A}\left(c^{2}+c^{{\dagger}2}\right)/2$ (2c)
$\hat{H}_{AB}$ $=g_{A}\left(a^{{\dagger}}b+ab^{{\dagger}}\right)+g_{A}\left(ab+a^{{\dagger}}b^{{\dagger}}\right)$ (2d)
$\hat{H}_{BC}$ $=g_{B}\left(b^{{\dagger}}c+bc^{{\dagger}}\right)+g_{B}\left(bc+b^{{\dagger}}c^{{\dagger}}\right)$ (2e)
The annihilation and creation operators ($a,b,c$) and ($a^{{\dagger}},b^{{\dagger}},c^{{\dagger}}$) satisfy the standard commutation relations: $\left[a,a^{{\dagger}}\right]=\mathbb{I};~{}\left[b,b^{{\dagger}}\right]=\mathbb{I};~{}\left[c,c^{{\dagger}}\right]=\mathbb{I}$. To study the system’s dynamics, we employ the covariance matrix formalism [40]. (See Appendix (A) for details.) First, we express Hamiltonian (1) as:
$\hat{H}=\mathbb{X}^{{\dagger}}H\mathbb{X}/2~{}~{}{\rm where}~{}~{}H=\begin{pmatrix}\bm{U}&\bm{V}\\ \bm{V}&\bm{U}\end{pmatrix}\,,$ (3)
$\hat{\mathbb{X}}=\left[a,b,c,a^{{\dagger}},b^{{\dagger}},c^{{\dagger}}\right]^{T}$ is the column vector of mode operators. The submatrices $\bm{U}$ and $\bm{V}$ are given by
$\bm{U}=\begin{pmatrix}\omega_{A}&g_{A}&0\\ g_{A}&\omega_{B}&g_{B}\\ 0&g_{B}&\omega_{A}\\ \end{pmatrix}\quad\bm{V}=\begin{pmatrix}\lambda_{A}&g_{A}&0\\ g_{A}&\lambda_{B}&g_{B}\\ 0&g_{B}&\lambda_{A}\end{pmatrix}$ (4)
The Hamiltonian in Eq. (1) is quadratic in the Bosonic creation and annihilation operators, guaranteeing that the time evolution preserves Gaussianity [41]. Consequently, if the system is initially in a Gaussian state, it remains Gaussian throughout [42]. We therefore restrict our analysis to Gaussian states ($\hat{\rho}_{g}$), which are fully characterized by their first and second moments of the quadrature operator vector $\hat{\mathbb{X}}$ and are defined as:
$\mathbf{d}_{i}=\langle\hat{\mathbb{X}}_{i}\rangle_{g};~{}~{}\bm{\sigma}_{nm}=\langle\{\hat{\mathbb{X}}_{n},\hat{\mathbb{X}}^{{\dagger}}_{m}\}\rangle{g}-2\langle\hat{\mathbb{X}}_{n}\rangle_{g}\langle\hat{\mathbb{X}}_{m}^{{\dagger}}\rangle_{g}$ (5)
[…]
The time evolution of the moments is governed by the symplectic transformation $S_{H}(t)=\exp(\Omega Ht)$, where the canonical symplectic form for this complex-operator basis is $\Omega=-i\,\text{diag}(1,1,1,-1,-1,-1)$. This transformation preserves the symplectic structure, $S_{H}(t)\,\Omega\,S_{H}^{T}(t)=\Omega$, thereby ensuring that the commutation relations remain invariant. The first and second moments evolve as:
$\mathbf{d}(t)=S_{H}(t)\,\mathbf{d}(0),\quad\bm{\sigma}(t)=S_{H}(t)\,\bm{\sigma}(0)\,S_{H}^{\dagger}(t).$ (6)
Since first moments do not affect quantum correlations such as entanglement or discord, we set $\mathbf{d}(0)=0$ without loss of generality and focus on the evolution of the covariance matrix $\bm{\sigma}(t)$.
To investigate the development of quantum correlations in this system, we study the time evolution of the covariance matrix $\bm{\sigma}(t)$ for the Hamiltonian (1). In realistic experimental scenarios — such as the proposed BMV-type setups — gravitationally mediated entanglement between two massive quantum systems (e.g., A and C) would be inferred from measurements performed over multiple experimental runs. As such experiments access only ensemble-averaged data, we evaluate long-time averages of ($\mathcal{E}_{\mathcal{N}}$) and ($\mathcal{D}$), computed over timescales much longer than the inverse normal-mode frequencies. These long-time averages qualitatively capture the entanglement structure mediated by oscillator B, reflecting what would be observed in ensemble measurements across many experimental runs.
III Time evolution of the Hamiltonian (1)
The time evolution of the covariance matrix $\bm{\sigma}(t)$ is obtained from the system’s normal mode frequencies $(k_{1},k_{2},k_{3})$ and the Bogoliubov coefficients $\bm{\alpha}$ and $\bm{\beta}$, which transforms $H$ into a diagonal form corresponding to the Williamson normal form [43], combined with the initial state’s covariance matrix (see Appendix (B) for details). To analyze quantum correlations between subsystems A and C, we trace out the mediator B from the time-evolved covariance matrix $\bm{\sigma}(t)$ to obtain the reduced state $\bm{\sigma}_{A-C}(t)$. The logarithmic negativity $\mathcal{E}_{\mathcal{N}}(t)$ and Gaussian quantum discord $\mathcal{D}(t)$ are then computed using standard techniques for Gaussian states [24, 25] (see Appendix (C) for details).
We proceed to analyze the system dynamics across a range of initial states and couplings, examining how quantum correlations depend on both system parameters and initial state preparation. This allows us to understand how the presence or absence of entanglement constrains the system and initial state parameters. As a starting point, we consider the most general separable Gaussian initial state: a product of single-mode squeezed states for each oscillator:
[…]
where each |$\ket{r,\theta}$ is a single-mode squeezed vacuum state:
|$\ket{r,\theta}=S(\zeta)\ket{0},\quad S(\zeta)=\exp{\left[(\zeta^{*}\hat{a}^{2}-\zeta\hat{a}^{\dagger 2})/2\right]}$ ) (8)
Here, $\zeta=re^{i\theta}$ is the complex squeezing parameter, with $r$ ($\theta$) as the squeezing strength (angle) [44]. While analytical solutions for $\mathcal{E}_{\mathcal{N}}(t)$ and $\mathcal{D}(t)$ are, in principle, possible for general Gaussian initial states, the complexity in asymmetric parameter regimes makes this approach cumbersome. Therefore, our primary method is numerical simulation to explore the full parameter space.
To gain physical insight, we perform a calculation in the weak coupling limit for the simplest configuration: an initial vacuum state ($r_{1}=r_{2}=r_{3}=0,\theta_{1}=\theta_{2}=\theta_{3}=0$), frequency-degenerate oscillators ($\omega_{A}=\omega_{B}=\omega$), symmetric couplings to the mediator ($g_{A}=g_{B}=g$), and no local squeezing interactions ($\lambda_{A}=\lambda_{B}=0$). In this regime, we consider the limit $\lambda_{p}\equiv\sqrt{2}g/\omega\ll 1$ limit. We find that the long-time averaged logarithmic negativity $\langle\mathcal{E}_{\mathcal{N}}\rangle$ between subsystems A and C takes the form:
$\langle\mathcal{E}_{\mathcal{N}}\rangle\approx\frac{\lambda_{p}}{\pi\ln 2}=\frac{\sqrt{2}}{\pi\ln 2}\frac{g}{\omega}$ (9)
(see Appendix (D) for details). Interestingly, in the weak coupling regime, the average entanglement varies inversely with the mediator frequency $\omega$, implying that lower-frequency mediators facilitate stronger effective correlations. However, the apparent divergence as $\omega\to 0$ is unphysical because the weak coupling condition ($\lambda_{p}\ll 1$) requires $g$ to vanish at least proportionally to $\omega$. This behavior is nonetheless reminiscent of entanglement enhancement seen in bipartite systems featuring zero modes [45, 46].
[…]
IV Numerical Analysis and physical regimes
To move beyond the analytically tractable weak-coupling regime, we numerically simulate the system’s dynamics using the full Bogoliubov coefficients ($\bm{\alpha},\bm{\beta}$) (See (B) for details). As mentioned above, our analysis focuses on two distinct physical regimes characterized by the mediator’s behavior, considering both initial squeezed and vacuum states. We present here the results for squeezed states, see Appendix (E) for vacuum initial state. The two limits delineate how the mediator’s dynamical accessibility and quantumness control entanglement generation across the system.
  1. 1.
    HMR: This regime is defined by the limit $\lambda_{B}\to\omega_{B}$, with $1-\lambda_{A}/\omega_{A}$ remaining finite. In this limit, the mediator oscillator B behaves as an effectively heavy, quasi-classical field mode. To understand its physical implications, consider the single-oscillator Hamiltonian with self-squeezing (cf. Eq. (2b)). When expressed in the position–momentum basis, after transformation, the Hamiltonian simplifies to:
    $\hat{H}_{0B}={\hat{p}_{B}^{2}}/{(2\tilde{m}_{B}})+\tilde{k}_{B}\,\hat{x}_{B}^{2}/2$ (10)
    Here, $\tilde{m}_{B}$ and $\tilde{k}_{B}$ are the renormalized parameters, given by:
    $\tilde{m}_{B}=\frac{m_{B}}{1-\lambda_{B}/\omega_{B}},~{}\tilde{\omega}_{B}^{2}=\omega_{B}^{2}-\lambda_{B}^{2},~{}\tilde{k}_{B}=\tilde{m}_{B}\tilde{\omega}_{B}^{2},$ (11)
    where $m_{B}$ is the bare mass of the mediator. In the critical limiting case $\lambda_{B}\to\omega_{B}$, the effective mass $\tilde{m}_{B}$ diverges ($\tilde{m}_{B}\to\infty$) while the renormalized frequency vanishes ($\tilde{\omega}_{B}\to 0$). Crucially, their product $\tilde{k}_{B}=\tilde{m}_{B}\tilde{\omega}_{B}^{2}$ remains finite and equal to twice the bare spring constant $m_{B}\omega_{B}^{2}$. Physically, this causes the kinetic term to vanish from the Hamiltonian (10), i.e.,
    $\hat{H}_{0B}\xrightarrow[\lambda_{B}\to\omega_{B}]{}\frac{1}{2}\tilde{k}_{B}\,\hat{x}_{B}^{2}.$ (12)
    This limit corresponds to an extreme self-interaction that effectively suppresses the canonical momentum $\hat{p}_{B}$, making $\hat{x}_{B}$ a conserved quantity. Hence, the oscillator no longer evolves dynamically but acts as a static background field fixed by its initial position $\hat{x}_{B}(0)$. This behavior is analogous to a cosmological constant [47], where vacuum energy is given by the constant scalar field potential, here represented by a fixed energy density parameterized by $\hat{x}_{B}(0)^{2}$.
    Thus, in HMR, the mediator loses its ability to coherently store or transmit quantum information. Its Hilbert space remains intact, but its dynamical role is effectively frozen: the canonical momentum $\hat{p}_{B}$ vanishes from the Hamiltonian, and the position $\hat{x}_{B}$ becomes a conserved quantity. This reduction mirrors constrained dynamics in gauge theories and systems with Hamiltonian constraints, where certain degrees of freedom become dynamically redundant or decouple from the evolution [48, 49].
    To visualize how entanglement between subsystems A and C depend on system parameters in both mediator regimes, we plot contour maps of the long-time averaged $\langle\mathcal{E}_{\mathcal{N}}\rangle$. (For $\langle\mathcal{D}\rangle$, see Appendix (E).) The first two plots on the left in Fig. (2) contain the plot of $\langle\mathcal{E}_{\mathcal{N}}\rangle$ and $\langle\mathcal{F}\rangle$ respectively as a function of $\omega_{A}$ and $g_{A}$ by fixing $\omega_{B}$, $\lambda_{A}$, $\lambda_{B}$, and $g_{B}$.
  2. 2.
    LMR: This regime is characterized by $\lambda_{A}\to\omega_{A}$, while $1-\lambda_{B}/\omega_{B}$ remains finite. In this scenario, the mediator B is not significantly self-squeezed, remaining highly quantum and susceptible to fluctuations. The last two plots on the right in Fig. (2) contain the plot of $\langle\mathcal{E}_{\mathcal{N}}\rangle$ and $\langle\mathcal{F}\rangle$ respectively as a function of $\omega_{B}$ and $g_{B}$ by fixing $\omega_{A}$, $\lambda_{A}$, $\lambda_{B}$, and $g_{A}$. (For $\langle\mathcal{D}\rangle$, see Appendix (E).)
V Results and Implications
Our numerical analysis, presented in Fig. 2 reveal that both (HMR and LMR) regimes can support generation of quantum correlations, quantified by $\langle\mathcal{E}_{\mathcal{N}}\rangle$ and $\langle\mathcal{D}\rangle$, between subsystems A and C. However, these regimes exhibit markedly different structural and spectral features in their entanglement.
The LMR consistently exhibits a smoother, more continuous entanglement landscape. Here, non-zero quantum correlations are broadly distributed across wider parameter ranges, indicating a robust ability of the “light” (dynamically active) mediator B to facilitate entanglement between A and C. This suggests that when the mediator can fluctuate significantly, it readily mediates quantum correlations.
In stark contrast, HMR presents a highly fragmented entanglement landscape. Entanglement emerges only within sharply localized islands in parameter space, surrounded by expansive regions where quantum correlations are strongly suppressed, often vanishing entirely. Despite the system’s integrability, the emergence of such sharply structured regions in the HMR is strikingly reminiscent of quantum many-body scars. In systems with scars, specific non-generic eigenstates exhibit unexpectedly low entanglement and resist thermalization, maintaining enhanced localization in Hilbert space. Our findings suggest an analogous form of scars, not necessarily in the many-body Hilbert space, but within the space of Gaussian states of this continuous-variable system. Further fidelity analysis support this: in the low-entanglement regions, the fidelity of the evolved state with the initial state remains remarkably close to unity over long times. This indicates dynamical stability and a resistance to spreading in the Hilbert space, akin to localized scarred states. Conversely, within the high-entanglement islands, fidelity oscillates significantly away from $1$, reflecting more complex, delocalized, and dynamically richer behavior. This coexistence of dynamically localized and delocalized states within an integrable system points towards selective resonance effects or subtle hidden symmetries that prevent full ergodic mixing in the continuous-variable phase space.
These differences are greatly reflected in the underlying eigenmode structure of the system, characterized by the normal mode frequencies $k_{1},k_{2},k_{3}$. In the HMR ($\lambda_{B}\to\omega_{B}$), one eigenmode approaches zero ($k_{2}\to 0$), signifying the “freezing” of the mediator’s kinetic degree of freedom, while the other two become degenerate ($k_{1}\to k_{3}$), implying a symmetric, coupled motion of A and C dominated by the static mediator. In the LMR ($\lambda_{A}\to\omega_{A}$), a different spectral signature emerges: two eigenmodes vanish simultaneously ($k_{1},k_{2}\to 0$), with the third remaining finite. This indicates a different type of effective decoupling or simplification of dynamics, possibly leaving the mediator’s own quantum fluctuations more dominant in mediating correlations. These contrasting signatures highlight a deeper asymmetry in the dynamical role of the mediator across the two regimes and may have implications for relativistic systems [18, 21].
Crucially, these results have significant implications for experimental tests of quantum gravity, such as BMV-type scenarios where two massive systems interact via a presumed quantum mediator. Our findings suggest that bipartite entanglement alone may offer only a partial diagnostic of the mediator’s quantum nature. Specifically, the absence of observed entanglement does not necessarily imply classical mediation. Instead, it may indicate that the mediator (analogous to oscillator B) is operating within a dynamically constrained subspace of its Hilbert space. This corresponds to the singular limit $\lambda_{B}\to\omega_{B}$ (HMR), wherein its ability to generate and transmit quantum correlations is effectively frozen, despite being fundamentally a quantum system. The observed vanishing entanglement in these specific parameter regions, even for non-zero quantum interactions, presents a critical caveat for entanglement-based protocols.
This singular limit, where the mediator’s kinetic energy is suppressed and its position becomes effectively static, may have a deeper physical interpretation within the context of quantum gravity. Given that gravity is inherently self-interacting and highly non-linear, such constrained dynamical behavior could be reminiscent of gravitational memory effects. In the context of BMV-type proposals, this suggests that even within a fully quantum framework of gravity, certain quasi-classical features (or strong self-interactions analogous to the self-squeezing term) could drive the mediating gravitational field into a subspace where it is dynamically inert and unable to coherently transmit quantum correlations. This implies that the quantum signature of gravity might be masked under specific, physically plausible conditions.
Therefore, for experimental tests aiming to definitively establish the quantum nature of gravity, it is paramount to account for such subtle dynamical constraints. Protocols must ensure that the hypothesized quantum mediator (e.g., gravitons or their effective field) operates in a regime where its quantum dynamical role is active and observable, rather than being effectively “frozen” into a quasi-classical configuration by strong self-interactions or specific environmental couplings. Our work provides a concrete, integrable model demonstrating how entanglement can vanish in a quantum system with non-zero interactions, serving as a critical benchmark for design and interpretation of future quantum gravity experiments.
Acknowledgements.
The authors are grateful to S. M. Chandran, Himadri Dhar, K. Hari and Y. Ibrahim for their valuable discussions and feedback on the earlier draft. The work is supported by SERB-CRG/2022/002348. The MHRD fellowship at IIT Bombay financially supports PGC.
Appendix A Symplectic formalism
This section provides a brief overview of the covariance matrix and symplectic formalism, which is particularly well-suited for systems governed by Hamiltonian that are quadratic in the canonical quadrature operators ($\hat{q}$ and $\hat{p}$), or equivalently in the Bosonic creation and annihilation operators ($\hat{a}$ and $\hat{a}^{\dagger}$). Instead of working in the phase-space picture via wavefunctions or characteristic functions of type “$s$” (denoted $\chi_{s}$), a more elegant and computationally efficient approach involves the use of the covariance matrix formalism, especially for Gaussian states. For Gaussian states, knowledge of the first and second moments is sufficient to fully reconstruct the state, as all higher-order moments can be expressed in terms of them due to Wick’s theorem [50].
We begin by defining a linear vector space (LVS) by collecting the annihilation and creation operators into a column vector of the form:
$\hat{\mathbb{X}}=\left[\hat{a}_{1},\hat{a}_{2},\ldots,\hat{a}_{n},\hat{a}^{{\dagger}}_{1},\hat{a}^{{\dagger}}_{2},\ldots,\hat{a}^{{\dagger}}_{n}\right]^{T}$
of length “$2n$”, where “$n$” represents the number of oscillator modes in the system. The canonical commutation relations between the various mode operators can now be recast as
$\left[\hat{\mathbb{X}}_{i},\hat{\mathbb{X}}^{{\dagger}}_{j}\right]=i\Omega_{ij}$ (13)
where, $\Omega$ is called the symplectic form given by,
$\Omega=-i\begin{pmatrix}\mathbb{I}_{n}&\mathbb{O}_{n}\\ \mathbb{O}_{n}&-\mathbb{I}_{n}\end{pmatrix}$ (14)
In the traditional density matrix formalism the time evolution of the system is given by a unitary operator $\hat{U}(t)$ such that $\rho_{0}\rightarrow\rho(t)=\hat{U}(t)\rho_{0}\hat{U}^{{\dagger}}(t)$, which in the covariance matrix formalism corresponds to a symplectic transformation of the first and second order moments. The first order moments transform as $\langle\hat{\mathbb{X}}_{0}\rangle\rightarrow\langle\hat{\mathbb{X}}_{0}\rangle(t)=S_{H}(t)\langle\hat{\mathbb{X}}_{0}\rangle$ while the second order moments in the form of the covariance matrix transform as [51]:
$\sigma_{0}\rightarrow\sigma(t)=S_{H}(t)\sigma_{0}S_{H}^{T}(t)\,,$
where the covariance matrix is defined as,
$\sigma_{ij}=\langle\{\hat{\mathbb{X}}_{i},\hat{\mathbb{X}}_{j}^{{\dagger}}\}\rangle-2\langle\hat{\mathbb{X}}_{i}\rangle\langle\hat{\mathbb{X}}^{{\dagger}}_{j}\rangle$ (15)
For time independent Hammitonian’s the symplectic evolution matrix is given by $S_{H}(t)=\exp[\Omega Ht]$ ($\hbar=1$). Where $H$ represents the Hamiltonian quadratic form, which appears when we write the total Hamiltonian in terms of the vector $\hat{\mathbb{X}}$ as $\hat{H}=\frac{1}{2}\hat{\mathbb{X}^{{\dagger}}}H\hat{\mathbb{X}}$. It is in general a hermitian matrix having the following structure,
$H=\begin{pmatrix}\bm{U}&\bm{V}\\ \bm{V}^{*}&\bm{U}^{*}\end{pmatrix}$ (16)
with $U=U^{{\dagger}}$ and $V=V^{T}$ due to hermicity of the Hamiltonian operator $\hat{H}$. This specific structure of $H$ arises due to the basis that we have chosen for our LVS, had we chosen a real basis for the column vector $\hat{\mathbb{X}_{r}}=[\hat{x}_{1},\hat{x}_{2},\ldots,\hat{x}_{n},\hat{p}_{1},\hat{p}_{2},\ldots,\hat{p}_{n}]^{T}$, then it is related to the column vector $\hat{\mathbb{X}}$ by
$\hat{\mathbb{X}}=\mathbb{U}\;\hat{\mathbb{X}_{r}}$ (17)
Where the basis changing matrix $\mathbb{U}$ elegantly reads
$\mathbb{U}=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{I}_{n}&i\mathbb{I}_{n}\\ \mathbb{I}_{n}&-i\mathbb{I}_{n}\end{pmatrix}$ (18)
Thus, any matrix $\mathbb{A}_{r}$ in the real basis will be written as,
$A=\mathbb{U}\;\mathbb{A}_{r}\;\mathbb{U}^{{\dagger}}$ (19)
in the complex basis. Thus if we write $\mathbb{A}_{r}$ in the form of block matrices,
$\mathbb{A}_{r}=\begin{pmatrix}\bm{L}&\bm{M}\\ \bm{N}&\bm{O}\end{pmatrix}$ (20)
The same matrix in the transformed basis takes the form,
$A=\begin{pmatrix}(\bm{L}+\bm{O})+i(\bm{N}-\bm{M})&(\bm{L}-\bm{O})+i(\bm{N}+\bm{M})\\ (\bm{L}-\bm{O})-i(\bm{N}-\bm{M})&(\bm{L}+\bm{O})-i(\bm{N}-\bm{M})\end{pmatrix}$
Thus, any matrix in the complex basis would have the same structure as that of Eq (16). For the case where $U$ and $V$ are real matrices, $H$ becomes a real symmetric positive definite (the antisymmetric component of $H$ only contributes as a constant to the Hamiltonian and hence ignored  [52]) $2n\times 2n$ matrix. The symplectic time evolution matrix $S_{H}$ satisfies $S_{H}\Omega S_{H}^{{\dagger}}=S_{H}^{{\dagger}}\Omega S_{H}=\Omega$, and in general can be written as,
$S_{H}(t)=\begin{pmatrix}\bm{A}(t)&\bm{B}(t)\\ \bm{B}^{*}(t)&\bm{A}^{*}(t)\end{pmatrix}$ (21)
Now according to Williamson’s theorem, any positive definite symmetric $2n\times 2n$ matrix can be diagonalised via a symplectic transformation and put in the Williamson normal form. Applying this theorem to $H$, we obtain
$H=S_{0}^{{\dagger}}\bm{K}S_{0}$ (22)
where $\bm{K}$ is a diagonal matrix of the form diag($k_{1},k_{2},\ldots k_{n},k_{1},k_{2},\ldots,k_{n}$) ($k_{i}\in\mathbb{R}^{+}$ for $i=1,2,\ldots,n$). The diagonal entries are called the symplectic eigenvalues of $H$, which are eigenvalues of the matrix $i\Omega H$. The general structure of the symplectic matrix $S_{0}$ is again of the same form as (16) and can be written as,
$S_{0}=\begin{pmatrix}\bm{\alpha}&\bm{\beta}\\ \bm{\beta}^{*}&\bm{\alpha}^{*}\end{pmatrix}$ (23)
Where the matrices $\bm{\alpha}$ and $\bm{\beta}$ are “$n\times n$”matrices called the Bogoliubov matrices. The symplectic condition $S_{0}\Omega S_{0}^{{\dagger}}=S_{0}^{{\dagger}}\Omega S_{0}=\Omega$ imposes further conditions on the Bogoliubov matrices called the Bogoliubov identities given as,
$\bm{\alpha}\bm{\alpha}^{{\dagger}}-\bm{\beta}\bm{\beta}^{{\dagger}}=\mathbb{I}_{2n}$ (24a)
$\bm{\alpha}\bm{\beta}^{{\dagger}}-\bm{\beta}\bm{\alpha}^{{\dagger}}=\mathbb{O}_{2n}$ (24b)
Using Eq. (22), we can simplify the symplectic time evolution matrix $S_{H}(t)$ as follows
$S_{H}(t)$ $=\exp[\Omega Ht]$
$=\exp[\Omega S_{0}^{{\dagger}}\bm{K}S_{0}t]$
$=\exp[S_{0}^{-1}S_{0}\Omega S_{0}^{{\dagger}}\bm{K}S_{0}t]$
$=\exp[S_{0}^{-1}\Omega\bm{K}S_{0}t]$
$=S_{0}^{-1}\exp[\Omega\bm{K}t]S_{0}$ (25)
However the symplectic form $\Omega$ is a diagonal matrix, hence using $S_{0}\Omega S_{0}^{{\dagger}}=\Omega$, we can rewrite Eq. (25) as
$S_{H}(t)$ $=-\Omega S_{0}^{{\dagger}}\Omega\exp[\Omega\bm{K}t]S_{0}$ (26)
Now if we introduce $\tilde{k}=$ diag($k_{1},k_{2},\ldots,k_{n}$) and expand the right hand side of Eq. (26) we obtain the equations,
$\bm{A}(t)$ $=\bm{\alpha}^{{\dagger}}e^{-i\tilde{k}t}\bm{\alpha}-\bm{\beta}^{T}e^{i\tilde{k}t}\bm{\beta}^{*}$ (27a)
$\bm{B}(t)$ $=\bm{\alpha}^{{\dagger}}e^{-i\tilde{k}t}\bm{\beta}-\bm{\beta}^{T}e^{i\tilde{k}t}\bm{\alpha}^{*}$ (27b)
With $\bm{A}^{T}(t)=\bm{A}(t)$ and $\bm{B}(t)=-\bm{B}^{{\dagger}}(t)$. Further using Eq. (16) and Eq. (22) we get the following equations,
$\bm{U}$ $=\bm{\alpha}^{{\dagger}}\tilde{k}\bm{\alpha}+\bm{\beta}^{T}\tilde{k}\bm{\beta}^{*}$ (28a)
$\bm{V}$ $=\bm{\alpha}^{{\dagger}}\tilde{k}\bm{\beta}+\bm{\beta}^{T}\tilde{k}\bm{\alpha}^{*}$ (28b)
Thus the entire system consists of solving the set of Eqs. (28) along with the constraint Eqs. (24), which gives the explicit expressions for $\bm{\alpha},\;\bm{\beta}$ and $\tilde{k}$. Once we obtain these quantities we can use Eqs. (27) to obtain $S_{H}(t)$ and then find $\langle\hat{\mathbb{X}_{0}}\rangle(t)$ and $\sigma(t)$. If at time “$t$”, the covariance matrix is given as ,
$\bm{\sigma}(t)=\begin{pmatrix}\bm{\mathcal{U}(t)}&\bm{\mathcal{V}(t)}\\ \bm{\mathcal{V}^{*}(t)}&\bm{\mathcal{U}^{*}(t)}\end{pmatrix}$ (29)
The explicit form for $\bm{\mathcal{U}}(t)$ and $\bm{\mathcal{V}}(t)$ can be obtained using the explicit form of $S_{H}(t)$ from Eq (21), which gives us,
$\bm{\mathcal{U}}(t)$ $=\bm{A}(t)\bm{\mathcal{U}}_{0}\bm{A}^{*}(t)+\bm{B}(t)\bm{\mathcal{V}}^{*}_{0}\bm{A}^{*}(t)$ (30)
$\quad-\bm{B}(t)\bm{\mathcal{U}}^{*}_{0}\bm{B}(t)-\bm{A}(t)\bm{\mathcal{V}}_{0}\bm{B}(t)$
$\bm{\mathcal{V}}(t)$ $=\bm{A}(t)\bm{\mathcal{V}}_{0}\bm{A}(t)+\bm{B}(t)\bm{\mathcal{U}}^{*}_{0}\bm{A}(t)$
$\quad-\bm{A}(t)\bm{\mathcal{U}}_{0}\bm{B}^{*}(t)-\bm{B}(t)\bm{\mathcal{V}}^{*}_{0}\bm{B}^{*}(t)$
Appendix B Solution to the general three Harmonic oscillator model
In this section, we present the explicit solution for the dynamics of the system discussed in main text [53]. The Hamiltonian quadratic form ($H$) of the system is composed of purely real parameters, as a result the entire covariance formalism can be reformulataed assuming all the matrices to be real. Thus the set of Eqs. (24) and Eqs. (24) and (28) reduce to
$\mathbb{I}_{2n}$ $=\bm{\alpha}\bm{\alpha}^{T}-\bm{\beta}\bm{\beta}^{T}$ (31a)
$\mathbb{O}_{2n}$ $=\bm{\alpha}\bm{\beta}^{T}-\bm{\beta}\bm{\alpha}^{T}$ (31b)
$\bm{U}$ $=\bm{\alpha}^{T}\tilde{k}\bm{\alpha}+\bm{\beta}^{T}\tilde{k}\bm{\beta}$ (31c)
$\bm{V}$ $=\bm{\alpha}^{T}\tilde{k}\bm{\beta}+\bm{\beta}^{T}\tilde{k}\bm{\alpha}$ (31d)
where $\bm{U}$ and $\bm{V}$ are given by Eq.(4) in the main text [53]. To check whether the system of equations at hand are consistent, we can count the number of independent unknowns and independent equations in them. Due to symmetry of the matrices $\bm{U}$ and $\bm{V}$, Eqs. (31c) - (31d) constitute 12 independent equations in 21 unknowns. However, the 21 unknowns are all not independent and are constrained by Eqs. (31a) – (31b) which are 9 in number. Thus, we have (21 - 9) = 12 unknowns and 12 independent equations.
We can recast these set of equations in a much more elegant way by introducing linear combinations of the Bogoliubov matrices as $\Delta_{+}=\bm{\alpha}+\bm{\beta}$ and $\Delta_{-}=\bm{\alpha}-\bm{\beta}$, then the Bogoliubov identies become a single matrix equation as $\Delta_{+}\Delta^{T}_{-}=\Delta_{-}\Delta^{T}_{+}=\mathbb{I}_{n}$ (where $n=3$ for this case), we can do further manipulation to write the equations for $\bm{U}$ and $\bm{V}$ in a better way by again considering linear combination of them as $\bm{U}_{+}=\bm{U}+\bm{V}$ and $\bm{U}_{-}=\bm{U}-\bm{V}$ and the corresponding equations become,
$\bm{U}_{+}=\Delta^{T}_{+}\tilde{k}\Delta_{+}$ (32a)
$\bm{U}_{-}=\Delta^{T}_{-}\tilde{k}\Delta_{-}$ (32b)
On utilizing the Bogoliubov identities we obtain much elegant looking equations
$\Delta_{-}\bm{U}_{+}=\tilde{k}\Delta_{+}$ (33a)
$\Delta_{+}\bm{U}_{-}=\tilde{k}\Delta_{-}$ (33b)
For the given choice of matrices $\bm{U}$ and $\bm{V}$, we can explicitly find our the symplectic eigenvalues of $H$ by writing the characterisitic polynomial of the matrix $i\Omega H$ under the assumptions $\omega_{1}>\lambda_{1},\omega_{2}>\lambda_{2}$, which gives
$f(\lambda)$ $=$ $(\lambda^{2}+\lambda_{1}^{2}-\omega_{1}^{2})(\lambda^{4}+R(\omega_{1}-\lambda_{1})(\omega_{2}-\lambda_{2})$ (34)
$+\lambda^{2}(\lambda_{1}^{2}+\lambda_{2}^{2}-\omega_{1}^{2}-\omega_{2}^{2}))$
$R$ $=$ $(\omega_{1}+\lambda_{1})(\omega_{2}+\lambda_{2})-4(g_{1}^{2}+g_{2}^{2})$
Solution to $f(\lambda)=0$ gives the symplectic eigenvalues,
$k_{1}$ $=\sqrt{\omega_{1}^{2}-\lambda_{1}^{2}}$ (35)
$k_{2}$ $=\frac{1}{\sqrt{2}}\sqrt{(\omega_{1}^{2}+\omega_{2}^{2}-\lambda_{1}^{2}-\lambda_{2}^{2})-\sqrt{Q}}$ (36)
$k_{3}$ $=\frac{1}{\sqrt{2}}\sqrt{(\omega_{1}^{2}+\omega_{2}^{2}-\lambda_{1}^{2}-\lambda_{2}^{2})+\sqrt{Q}}$ (37)
$Q$ $=(\omega_{1}^{2}+\omega_{2}^{2}-\lambda_{1}^{2}-\lambda_{2}^{2})^{2}-$
$4(\omega_{1}-\lambda_{1})(\omega_{2}-\lambda_{2})[(\omega_{1}+\lambda_{1})(\omega_{2}+\lambda_{2})-4(g_{1}^{2}+g_{2}^{2})]$
For the symplectic eigenvalues to be real positive, we require $Q\geq 0$. This imposes the following constraint on the system parameters,
$g_{1}^{2}+g_{2}^{2}\leq\frac{1}{4}(\omega_{1}+\lambda_{1})(\omega_{2}+\lambda_{2})$ (38)
With the symplectic eigenvalues already at hand, we can go ahead to solve the set of Eqs. (28), however, now we have to eliminate the three redundant equations and only solve for the 9 unknowns from 9 equations. Now using these equations, we can write down the following equations,
$\beta_{11}$ $=\frac{\alpha_{11}(k_{1}-\omega_{1}+\lambda_{1})}{(k_{1}+\omega_{1}-\lambda_{1})}$ (39)
$\beta_{12}$ $=\frac{\alpha_{12}(k_{1}-\omega_{2}+\lambda_{2})}{(k_{1}+\omega_{2}-\lambda_{2})}$ (40)
$\beta_{13}$ $=\frac{\alpha_{13}(k_{1}-\omega_{1}+\lambda_{1})}{(k_{1}+\omega_{1}-\lambda_{1})}$ (41)
$\beta_{21}$ $=\frac{\alpha_{21}(k_{2}-\omega_{1}+\lambda_{1})}{(k_{2}+\omega_{1}-\lambda_{1})}$ (42)
$\beta_{22}$ $=\frac{\alpha_{22}(k_{2}-\omega_{2}+\lambda_{2})}{(k_{2}+\omega_{2}-\lambda_{2})}$ (43)
$\beta_{23}$ $=\frac{\alpha_{23}(k_{2}-\omega_{1}+\lambda_{1})}{(k_{2}+\omega_{1}-\lambda_{1})}$ (44)
$\beta_{31}$ $=\frac{\alpha_{31}(k_{3}-\omega_{1}+\lambda_{1})}{(k_{3}+\omega_{1}-\lambda_{1})}$ (45)
$\beta_{32}$ $=\frac{\alpha_{32}(k_{3}-\omega_{2}+\lambda_{2})}{(k_{3}+\omega_{3}-\lambda_{2})}$ (46)
$\beta_{33}$ $=\frac{\alpha_{33}(k_{3}-\omega_{1}+\lambda_{1})}{(k_{3}+\omega_{1}-\lambda_{1})}$ (47)
Thus, we find the $\bm{\beta}$ matrix coefficients to be proportional to the corresponding coefficients of the $\bm{\alpha}$ matrix. Now we can use the above in the Bogoliubov identities to eliminate all the $\bm{\beta}$ coefficients,
$(\alpha_{11}^{2}+\alpha_{13}^{2})\frac{4k_{1}(\omega_{1}-\lambda_{1})}{(k_{1}+\omega_{1}-\lambda_{1})^{2}}\;+\alpha_{12}^{2}\frac{4k_{1}(\omega_{2}-\lambda_{2})}{(k_{1}+\omega_{2}-\lambda_{2})^{2}}=1$ (48)
$(\alpha_{21}^{2}+\alpha_{23}^{2})\frac{4k_{2}(\omega_{1}-\lambda_{1})}{(k_{2}+\omega_{1}-\lambda_{1})^{2}}\;+\alpha_{22}^{2}\frac{4k_{2}(\omega_{2}-\lambda_{2})}{(k_{2}+\omega_{2}-\lambda_{2})^{2}}=1$ (49)
$(\alpha_{31}^{2}+\alpha_{33}^{2})\frac{4k_{3}(\omega_{1}-\lambda_{1})}{(k_{3}+\omega_{1}-\lambda_{1})^{2}}\;+\alpha_{32}^{2}\frac{4k_{3}(\omega_{2}-\lambda_{2})}{(k_{3}+\omega_{2}-\lambda_{2})^{2}}=1$ (50)
The above equations arise from the diagonal constraints obtained by substituting the $\bm{\beta}$ matrix coefficients in Eq. (31a). The off-diagonal constraints give the following:
$\frac{(\alpha_{11}\alpha_{21}+\alpha_{13}\alpha_{23})(\omega_{1}-\lambda_{1})}{(k_{1}+\omega_{1}-\lambda_{1})(k_{2}+\omega_{1}-\lambda_{1})}\;+\frac{\alpha_{12}\alpha_{22}(\omega_{2}-\lambda_{2})}{(k_{1}+\omega_{2}-\lambda_{2})(k_{2}+\omega_{2}-\lambda_{2})}=0$ (51)
$\frac{(\alpha_{11}\alpha_{31}+\alpha_{13}\alpha_{33})(\omega_{1}-\lambda_{1})}{(k_{1}+\omega_{1}-\lambda_{1})(k_{3}+\omega_{1}-\lambda_{1})}\;+\frac{\alpha_{12}\alpha_{32}(\omega_{2}-\lambda_{2})}{(k_{1}+\omega_{2}-\lambda_{2})(k_{3}+\omega_{2}-\lambda_{2})}=0$ (52)
$\frac{(\alpha_{21}\alpha_{31}+\alpha_{23}\alpha_{33})(\omega_{1}-\lambda_{1})}{(k_{2}+\omega_{1}-\lambda_{1})(k_{3}+\omega_{1}-\lambda_{1})}\;+\frac{\alpha_{22}\alpha_{32}(\omega_{2}-\lambda_{2})}{(k_{2}+\omega_{2}-\lambda_{2})(k_{3}+\omega_{2}-\lambda_{2})}=0$ (53)
There are three more equations from the second Bogoliubov identities, however they turn out to be exactly the same as the above equations, as already expected since we obtained the symplectic eigenvalues independent of these equations. Now using the above equations and after a really lengthy algebra, we obtain the following for the $\alpha$ and $\beta$ coefficients,
Table 1: Expressions for the Bogoliubov coefficients $\alpha_{ij}$ and $\beta_{ij}$.
$j=1$ $j=2$ $j=3$
$\alpha_{1j}$ $\dfrac{(k_{1}+\omega_{1}-\lambda_{1})x_{1}}{2\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$ 0 $-\dfrac{(k_{1}+\omega_{1}-\lambda_{1})y_{1}}{2\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$
$\alpha_{2j}$ $\dfrac{(k_{2}+\omega_{1}-\lambda_{1})y_{1}y_{2}}{2\sqrt{k_{2}(\omega_{1}-\lambda_{1})}}$ $\dfrac{(k_{2}+\omega_{2}-\lambda_{2})x_{2}}{2\sqrt{k_{2}(\omega_{2}-\lambda_{2})}}$ $\dfrac{(k_{2}+\omega_{1}-\lambda_{1})x_{1}y_{2}}{2\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$
$\alpha_{3j}$ $-\dfrac{(k_{3}+\omega_{1}-\lambda_{1})y_{1}x_{2}}{2\sqrt{k_{3}(\omega_{1}-\lambda_{1})}}$ $\dfrac{(k_{2}+\omega_{2}-\lambda_{2})y_{2}}{2\sqrt{k_{3}(\omega_{2}-\lambda_{2})}}$ $-\dfrac{(k_{3}+\omega_{1}-\lambda_{1})x_{1}x_{2}}{2\sqrt{k_{3}(\omega_{1}-\lambda_{1})}}$
$\beta_{1j}$ $\dfrac{(k_{1}-\omega_{1}+\lambda_{1})x_{1}}{2\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$ 0 $-\dfrac{(k_{1}-\omega_{1}+\lambda_{1})y_{1}}{2\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$
$\beta_{2j}$ $\dfrac{(k_{2}-\omega_{1}+\lambda_{1})y_{1}y_{2}}{2\sqrt{k_{2}(\omega_{1}-\lambda_{1})}}$ $\dfrac{(k_{2}-\omega_{2}+\lambda_{2})x_{2}}{2\sqrt{k_{2}(\omega_{2}-\lambda_{2})}}$ $\dfrac{(k_{2}-\omega_{1}+\lambda_{1})x_{1}y_{2}}{2\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$
$\beta_{3j}$ $-\dfrac{(k_{3}-\omega_{1}+\lambda_{1})y_{1}x_{2}}{2\sqrt{k_{3}(\omega_{1}-\lambda_{1})}}$ $\dfrac{(k_{2}-\omega_{2}+\lambda_{2})y_{2}}{2\sqrt{k_{3}(\omega_{2}-\lambda_{2})}}$ $-\dfrac{(k_{3}-\omega_{1}+\lambda_{1})x_{1}x_{2}}{2\sqrt{k_{3}(\omega_{1}-\lambda_{1})}}$
Where the phase angles $\theta_{1}$, $\theta_{2}$ are defined as,
$x_{1}$ $=\cos{\theta_{1}}=\frac{g_{1}^{2}}{\sqrt{g_{1}^{2}+g_{2}^{2}}}$ (54)
$y_{1}$ $=\sin{\theta_{1}}=\frac{g_{2}^{2}}{\sqrt{g_{1}^{2}+g_{2}^{2}}}$ (55)
$x_{2}$ $=\cos{\theta_{2}}=\frac{(k_{2}^{2}+\lambda_{1}^{2}-\omega_{1}^{2})}{\sqrt{T}}$ (56)
$y_{2}$ $=\sin{\theta_{2}}=\frac{2\sqrt{(\omega_{1}-\lambda_{1})(\omega_{2}-\lambda_{2})(g_{1}^{2}+g_{2}^{2})}}{\sqrt{T}}$ (57)
$T$ $=(k_{2}^{2}+\lambda_{1}^{2}-\omega_{1}^{2})^{2}+4(\omega_{1}-\lambda_{1})(\omega_{2}-\lambda_{2})(g_{1}^{2}+g_{2}^{2})$
With the above Bogoliubov coefficients and the eigenmode frequencies, we can use the set of Eqs. (27) to solve for the dymanics of the system for any given initial state.
Appendix C Quantum Measures
To quantify quantum correlations between subsystems A and C, we compute the logarithmic negativity $\mathcal{E}_{\mathcal{N}}(t)$ and quantum discord $\mathcal{D}(t)$, both derived from the reduced covariance matrix $\bm{\sigma}_{A-C}(t)$, obtained by tracing out the mediator B.
For Gaussian states, entanglement is evaluated via the partial transpose operation, implemented by applying the matrix $\Lambda$ [54], yielding the partially transposed covariance matrix:
$\tilde{\bm{\sigma}}_{A-C}(t)=\Lambda\,\bm{\sigma}_{A-C}(t)\,\Lambda.$
The logarithmic negativity is computed from the symplectic eigenvalues $\tilde{\nu}_{i}(t)$ of $\tilde{\bm{\sigma}}_{A-C}(t)$, which are the eigenvalues of $i\Omega^{\prime}\tilde{\bm{\sigma}}_{A-C}(t)$, where $\Omega^{\prime}$ is the reduced symplectic form [24]. Explicitly,
$\mathcal{E}_{\mathcal{N}}(t)=\text{max}\left[0,-\sum_{i:\tilde{\nu_{i}(t)<1}}\log_{2}(\tilde{\nu}_{i}(t))\right]$ (58)
Quantum discord $\mathcal{D}(t)$ is computed from $\bm{\sigma}_{A-C}(t)$ :
$\mathcal{D}(t)=$ $f\left(\sqrt{B(t)}\right)-f\left(\nu_{-}(t)\right)-f\left(\nu_{+}(t)\right)$
$+$ $\inf_{\bm{\sigma}_{0}}f\left(\sqrt{\det\epsilon}\right)$ (59)
where $f(x)=\frac{x+1}{2}\log_{2}\left(\frac{x+1}{2}\right)-\frac{x-1}{2}\log_{2}\left(\frac{x-1}{2}\right)$, $B(t)$ is the determinant of the correlation block in $\bm{\sigma}_{A-C}(t)$, and $\nu_{\pm}(t)$ are the symplectic eigenvalues of $\bm{\sigma}_{A-C}(t)$, computed as the eigenvalues of $i\Omega^{\prime}\bm{\sigma}_{A-C}(t)$. Note that while the symplectic spectrum of the partially transposed covariance matrix may contain negative entries due to the mathematical diagonalization procedure, the physical symplectic eigenvalues relevant for computing the logarithmic negativity are defined as the absolute values of these quantities. The final term involves minimizing the determinant of the conditional covariance matrix over all single-mode Gaussian states $\bm{\sigma}_{0}$ and is defined as
$\epsilon=\bm{\sigma}_{A}-\bm{\sigma}_{B}(\bm{\sigma}_{C}+\bm{\sigma}_{0})^{-1}\bm{\sigma}_{B}^{\mathrm{T}}$ (60)
where $\bm{\sigma}_{A}$, $\bm{\sigma}_{B}$, and $\bm{\sigma}_{C}$ are the block matrices of the reduced covariance matrix $\bm{\sigma}_{A-C}(t)$:
$\bm{\sigma}_{A-C}(t)=\begin{pmatrix}\bm{\sigma}_{A}&\bm{\sigma}_{B}\\ \bm{\sigma}_{B}^{\mathrm{T}}&\bm{\sigma}_{C}\end{pmatrix},$ (61)
and $\bm{\sigma}_{0}$ corresponds to the covariance matrix of a single-mode Gaussian measurement on subsystem C. The minimum value of the determinant of $\epsilon$ is given in Eq. (67), where $\mathcal{A}=\det(\bm{\sigma}_{A})$, $\mathcal{B}=\det(\bm{\sigma}_{B})$, $\mathcal{C}=\det(\bm{\sigma}_{C})$, and $\mathcal{D}=\det(\bm{\sigma}_{A-C})$. The fidelity between two covariance matrices $V_{1}$ and $V_{2}$ is defined as [31]:
$F_{0}(V_{1},V_{2})$ $=\frac{F_{\text{tot}}}{\sqrt[4]{\det(V_{1}+V_{2})}}$ (62)
$F_{\text{tot}}^{4}$ $=\det\left[2\left(\sqrt{\mathbb{I}+\frac{(V_{\text{aux}}\Omega)^{-2}}{4}}+\mathbb{I}\right)V_{\text{aux}}\right]$ (63)
$=\det\left[\left(\sqrt{\mathbb{I}-W_{\text{aux}}^{-2}}+\mathbb{I}\right)W_{\text{aux}}\,i\Omega\right]$ (64)
where the auxiliary matrices $V_{\text{aux}}$ and $W_{\text{aux}}$ are given by
$V_{\text{aux}}$ $:=\Omega^{T}(V_{1}+V_{2})^{-1}\left(\frac{\Omega}{4}+V_{2}\Omega V_{1}\right)$ (65)
$W_{\text{aux}}$ $:=-2V_{\text{aux}}i\Omega=-(W_{1}+W_{2})^{-1}(\mathbb{I}+W_{2}W_{1})$ (66)
Here, $\Omega$ is the symplectic form, and $W_{i}:=-2V_{i}i\Omega$.
$E^{\text{min}}=\inf_{\bm{\sigma}_{0}}\det(\epsilon)=\begin{cases}\displaystyle\frac{2\mathcal{C}^{2}+(-1+\mathcal{B})(-\mathcal{A}+\mathcal{D})+2|\mathcal{C}|\sqrt{\mathcal{C}^{2}+(-1+\mathcal{B})(-\mathcal{A}+\mathcal{D})}}{(-1+\mathcal{B})^{2}}&\text{if }(\mathcal{D}-\mathcal{A}\mathcal{B})^{2}\leq(1+\mathcal{B})\mathcal{C}^{2}(\mathcal{A}+\mathcal{D})\\ \displaystyle\frac{\mathcal{A}\mathcal{B}-\mathcal{C}^{2}+\mathcal{D}-\sqrt{\mathcal{C}^{4}+(-\mathcal{A}\mathcal{B}+\mathcal{D})^{2}-2\mathcal{C}^{2}(\mathcal{A}\mathcal{B}+\mathcal{D})}}{2\mathcal{B}}&\text{otherwise}\end{cases}$ (67)
Appendix D Log Negativity in the weak coupling limit
In this section, we provide the details on the calculation of $\langle\mathcal{E}_{\mathcal{N}}\rangle$ in the weak-coupling regime with symmetric oscillators with $\omega_{A}=\omega_{B}=\omega$, $\lambda_{A}=\lambda_{B}=0$, $g_{A}=g_{B}=g$ and we define the dimensionless coupling parameter $\lambda_{p}=\sqrt{2}g/\omega$. In this limit, we perform a perturbative expansion of the eigenmode frequencies $k_{1},k_{2},k_{3}$ [Eqs. (35)–(37)] and the Bogoliubov coefficients (Table 1) in powers of $\lambda_{p}$, retaining only terms up to linear order and discarding higher-order corrections.
$\bm{\sigma}_{A-C}(t)\approx\mathbb{I}_{4}+\lambda_{p}\Delta+\mathcal{O}({\lambda_{p}^{2}})$ (68)
where $\mathbb{I}_{4}$ is $4\times 4$ identity matrix and $\Delta$ is defined as,
$\Delta=-\sin{(2\omega\lambda_{p}t)}R\begin{pmatrix}\mathbb{I}_{2}&\mathbb{I}_{2}\\ \mathbb{I}_{2}&\mathbb{I}_{2}\end{pmatrix}$ )
[…]
with $\mathbb{I}_{2}$ being the $2\times 2$ identity matrix. The symplectic eigenvalues of the partial transposed reduced covariance matrix $\tilde{\bm{\sigma}}_{A-C}(t)$ then becomes,
$\tilde{\nu}_{1+}$$\approx$ $\sqrt{1+\lambda_{p}\sin{(2\omega\lambda_{p}t)}}$ (70a)
$\tilde{\nu}_{1-}$$\approx$ $-\sqrt{1+\lambda_{p}\sin{(2\omega\lambda_{p}t)}}$ (70b)
$\tilde{\nu}_{2+}$$\approx$ $\sqrt{1-\lambda_{p}\sin{(2\omega\lambda_{p}t)}}$ (70c)
$\tilde{\nu}_{2-}$$\approx$ $-\sqrt{1-\lambda_{p}\sin{(2\omega\lambda_{p}t)}}$ (70d)
Using Eqs. (70) in Eq. (58), we obtain,
$\mathcal{E}_{\mathcal{N}}(t)\approx\frac{\lambda_{p}|\sin{(2\sqrt{2}gt)}|}{2\ln{(2)}}+\mathcal{O}({\lambda_{p}^{2}})$ ) (71)
Thus, the average value of $\mathcal{E}_{\mathcal{N}}(t)$ in the weak-coupling regime becomes,
$\langle\mathcal{E}_{\mathcal{N}}\rangle=\frac{\lambda_{p}}{\pi\ln 2}=\frac{\sqrt{2}}{\pi\ln 2}\frac{g}{\omega}$ (72)
Figure 3: Contour plots of time-averaged logarithmic negativity $\langle\mathcal{E}_{\mathcal{N}}\rangle$ and Fidelity $\langle\mathcal{F}\rangle$. (a, b)(HMR) use the parameter set $\omega_{B}=50$, $\lambda_{B}=49.9$, $\lambda_{A}=0.001$, $g_{B}=2.0$, (c, d)(LMR) use $\omega_{A}=5.0$, $\lambda_{A}=4.999$, $\lambda_{B}=\omega_{B}/1.5$, $g_{B}=1.4$
(a) HMR ($\omega_{B}=50.0$, $\lambda_{B}=49.9$, $\lambda_{A}=0.001$, $g_{B}=2.0$)
(b) HMR ($\omega_{B}=5.0$, $\lambda_{B}=4.9$, $\lambda_{A}=0.0001$, $g_{B}=1.0$)
(c) LMR ($\omega_{A}=5.0$, $\lambda_{A}=4.999$, $\lambda_{B}=\omega_{B}/1.5$, $g_{B}=1.4$)
(d) LMR ($\omega_{A}=5.0$, $\lambda_{A}=4.999$, $\lambda_{B}=\omega_{B}/1.5$, $g_{B}=1.4$)
Figure 4: Contour plots of time-averaged quantum discord $\langle\mathcal{D}\rangle$ in two contrasting mediator regimes. (a, c) Vacuum initial state, corresponding to zero squeezing: $r_{1}=r_{2}=r_{3}=0$, $\theta_{1}=\theta_{2}=\theta_{3}=0$; (b, d) Squeezed initial state with $r_{1}=(0.1,0.5)$, $r_{2}=(1.0,2.0)$, $r_{3}=(0.1,0.5)$ and squeezing angles $\theta_{1}=(-\pi/3,0)$, $\theta_{2}=(\pi/4,0)$, $\theta_{3}=(\pi/6,0)$.
Appendix E Plots for Vacuum Initial state
In Fig. (3) we provide the contour maps for $\langle\mathcal{E}_{\mathcal{N}}\rangle$, and $\langle\mathcal{F}\rangle$ for the case where the initial state of the system is vacuum. For the case of initial squeezed states, the plots were provided in Fig. (2) in the main text [53]. In Fig. (4) above we provide the countour plots for quantum discord $\langle\mathcal{D}\rangle$ both for initial squeezed state and initial vacuum state. Note that the behaviour of quantum discord resembles that of logarithmic negativity.